OpenStudy (anonymous):

Find an equation for the derivative of the given function: y=2^x lnx The answer should be -sin(lnx)(1/x). Can someone show me how to get there? Thanks!!

OpenStudy (anonymous):

use the product rule f'(x) = 2^x * 1/x + lnx * f'(2*x) let y = 2^x lny = x ln2 1/y dy/dx = ln 2 dy.dx = 2^x * ln2 f'(x) = 2^x * 1/x + lnx* 2*x* ln x = 2^x (1/x + (lnx )^2) - not quite what you have ! i might have missed something here

OpenStudy (anonymous):

I'm sorry, I wrote down the answer to the wrong problem. The answer should be 2^xlnx[ln2+1/(xlnx)]

OpenStudy (anonymous):

Or if this is more clear: (2^x)lnx[ln2+1/(xlnx)]

OpenStudy (anonymous):

oh i made an error in the last step its f'(x) = 2^x * 1/x + ln2 * 2^x * ln x = 2^x 2^x ln2 lnx --- + x = 2^x + 2^x ln2 x lnx ----------------- x = 2^x ( 1 + ln2 x lnx ) ----------------- x

OpenStudy (anonymous):

- cant seem to get your answer - i'd better try wolfram alpha

OpenStudy (anonymous):

Thank you so much for your help! I'll try to figure the rest out myself! Thank you!!!

OpenStudy (anonymous):

wolfram gives same answer as me your welcome